OpenStudy (anonymous):

Find an equation for the tangent line to the graph of f(x) = arctanx at the point where x = 1

OpenStudy (anonymous):

Find the slope of the curve at the point first.

OpenStudy (anonymous):

Would it be \[1/\sqrt{2}\]

OpenStudy (abb0t):

find \(f'(x)\)

OpenStudy (abb0t):

evaluated at x = 1

OpenStudy (anonymous):

That's what I attempted to do. lol. I tried it again, and using the formula from the textbook I got 0 this time. ><

OpenStudy (abb0t):

\(\frac{d}{dx} [arctan(x)] = \frac{1}{x^2+1}\)

OpenStudy (anonymous):

Oh! So m = 1?

OpenStudy (anonymous):

1/2 I mean!

OpenStudy (abb0t):

\[\frac{1}{x^2+1}|_{\left| _{x=1} \right|} = \frac{1}{2}\]

OpenStudy (anonymous):

So y = (1/2) * 1 + b. What would y be?

OpenStudy (abb0t):

why would u multiply 1+b? Where did that even come from?!

OpenStudy (anonymous):

y = mx + b. I just plugged in m and x. O.O

OpenStudy (abb0t):

well, \(\frac{1}{2}\) is your slope, so yes, m.

OpenStudy (anonymous):

And x = 1, right? The only element I seem to need is y, then I could solve for b and give the equation for the tangent line at x = 1.

OpenStudy (abb0t):

yes

OpenStudy (anonymous):

Alright, so arctan(1) comes out to 45. So it would then be 45 = 1/2 +b --> 44.5 = b --> y = (1/2)x + 44.5. ... I think. xD

OpenStudy (abb0t):

No. No. No. y = \(\frac{1}{2}x+b\)

OpenStudy (anonymous):

Why not give the b value?